Solve for n
2n+3-(4n+6)=9

Answer by ptaylor(2198)

Solve for n
2n+3-(4n+6)=9 clear parens:
2n+3-4n-6=9 collect like terms:
-2n-3=9 add 3 to both sides:
-2n-3+3=9+3
-2n=12
n=-6---------------------------------answer

ck
2(-6)+3-(4(-6)+6)=9
-12+3+24-6=9
9=9

Hope this helps---ptaylor